A group of friends takes a day-long tubing trip down a river. The company that offers the tubing trip charges $15 to rent a tube for a person to use and $7.50 to rent a "cooler" tube, which is used to carry food and water in a cooler. The friends spend $360 to rent a total of 26 tubes. How many of each type of tube do they rent?
Answer by checkley79(3341)

P+C=26
P=26-C
15P+7.50C=360
15(26-C)+7.50C=360
390-15C+7.50C=360
-7.50C=360-390
-7.5C=-30
C=30/7.50
C=4 COOLER RAFTS WERE RENTED.
P+4=26
P=26-4
P=22 PEOPLE RAFTS WERE RENTED.
PROOF:
15*22+7.50*4=360
330+30=360
360=360